Summary


Cross-disperser won't stop moving

Symptom
When a cross disperser move is requested, the XDANGL keyword starts changing and never stops; it cycles continuously within a certain range of values.
Problem
The cross disperser motor is confused.
Solution
  1. From the window manager's pulldown menu, select the HIRES Control Menu option, then release on the item labeled Stop Cross-disperser Oscillation.
  2. Alternatively, type the following command on lehoula:
    stop_cd_drift
  3. If this does not halt the cross-disperser, contact your support astronomer.

Shutter appears to fail

Symptom
During a non-dark exposure, the color-coded box on the exposure control GUI stays black indicating a closed shutter.
Problem
The shutter is opening but not enough to trigger the limit switch. The EXPTIME keyword in the fits header will also be fractions of a second rather than the desired exposure time. This may occur during humid conditions or when the shutter is failing.
Solution
  • No solution needed, unless the exposure level implies that the shutter really was closed. Shutter replacement may be necessary.

  • Rotator is not responding

    Symptom
    The rotator will not initialize.
    Problem
    The rotator daemon needs to be restarted or it is running as another user.
    Solution
  • Use the ct command to see what daemons are running (see example output below).
  • Try executing hires restart watch_hirot
  • Example ct output:

    [21] hires1@lehoula: ct

    ============================================================================
    Checking HIRES tasks...

    MANDATORY daemons on LEHOULA (owner 'hireseng'; 6 tasks):
    traffic, infoman, watch_ccd, watch_expo, dispatcher2, lickserv2

    OBSERVER processes on LEHOULA (after a full startup; owner 'YOU'; 4 tasks):
    write_image, obsr, ds9, ds9relay
    ============================================================================

    ...Checking LEHOULA for necessary daemon tasks ...
    OWNER PID TIME PROCESS
    hireseng 9362 19:10:24 0:05 /local/kroot/bin/watch_hirot dcs
    hireseng 29932 Aug 03 6:44 lickserv2 -s hiccd -port 5686 -wname h_watch_ccd -lsname h_lickserv2 -
    hireseng 29800 Aug 03 15:11 traffic
    hires1 14514 19:31:15 0:00 dispatcher2 -s expo -n 0 -l dy
    hireseng 9302 19:10:21 0:00 /bin/csh /local/kroot/bin/watch_hirot_monitor
    hireseng 9330 19:10:24 0:00 /bin/csh /local/kroot/bin/watch_hirot_rotinit
    hireseng 9332 19:10:24 0:00 /bin/csh /local/kroot/bin/watch_hirot_rotstby
    hireseng 9334 19:10:24 0:00 /bin/csh /local/kroot/bin/watch_hirot_rothalt
    hires1 14606 19:31:20 0:00 /local/kroot/bin/watch_expo
    hireseng 28169 Jul 25 8:14 watch_ccd /local/kroot/data/music/info.hiccd/header_info hiccd hires e
    hireseng 29833 Aug 03 0:01 infoman h_infom v=hiccd
    hireseng 29846 Aug 03 0:31 watch_ccd /local/kroot/data/music/info.hiccd/header_info hiccd hires e

    ...Checking LEHOULA for observer tasks ...
    OWNER PID TIME PROCESS
    hireseng 15483 19:34:12 0:26 /local/kroot/bin/obsr -2l /local/k
    hires1 14921 19:32:12 0:22 /local/kroot/bin/obsr -2l /local/k
    hireseng 21712 Aug 14 12:22 /local/kroot/bin/obsr -2l /local/k
    hires1 14886 19:32:11 0:00 write_image -s hiccd -wcname h_watch_ccd -lsname h_lickserv2 -winame h
    hires1 14898 19:32:12 0:10 /local/kroot/bin/obsr -2lv /local/
    hires1 14963 19:32:44 0:00 /local/ucolick/NewDS9/3.0.2/saods9/lib/ucolick/ds9relay

    Exposure Meter Errors

    Symptom
    The exposure meter is not responding to commands or error windows are popping up (e.g. "Heartbeat expo. CTRL0CLK has stopped").
    Problem
    The exposure meter is in an unhappy state and needs to be restarted.
    Solution
  • Attempt to restart the watch_expo daemon: hires restart watch_expo
  • If that fails with an error (e.g. Error: Can't read Exposure meter dispatcher status: mread select timeout.), then do an "End of Night Shutdown" and then restart all the GUIs with the "Start ALL HIRES GUIs" from the background menu.

  • Guider Bias Mismatch

    Symptom
    The two channels of the guider are very different (15k on left, 9k on right)
    Problem
    The camera biases have been set to 0 instead of default levels (515, 460) or so
    Solution
  • In MAGIQ OA UI, open "Extended Camera Controls"
  • Click "Load Defaults"